Cindy is an 11-year-old, 79 lb child who sliced her forearm open on a sharp piece of metal while helping her father clean out the garage. At the emergency room, the wound was cleaned up and the physician determined that she would need several sutures. Cindy is very upset, and the physician orders Midazolam for injection to sedate her and help decrease her anxiety. However, when Cindy sees the needle, she becomes very agitated, so the physician changes the order to Midazolam HCl syrup 18 mg PO. On hand, the emergency room has a 118 mL bottle of Midazolam HCl syrup 2mgmL 2 ⁢   m g m L . According to the package insert, the usual oral dose is 0.25 to 1 mg/kg. Determine if the ordered dose is safe. If it is safe, determine the amount to administer: _______________. If the dose is not safe, write: not safe. mL.
Answer #1

ANSWER.

* Calculation the amount of drug is safe or not.

             Usual oral dose = 0.25 mg to 1 mg/kg.

            Weight of the patient = 79 lb.

      Convert lb to kg for calculation

                 1 lb = 0.454 kg

        Weight of the patient = 79 * 0.454

                                          = 35.9 kg.

      Dose range for this patient = 0.25 * 35.9 to 1 * 35.9

                                                     = 8.9 mg to 35.9 mg.

                    Ordered dose = 18 mg.

* So ordered dose is safe for this patient.

* Calculation of amount (ml) of drug need to administer.

           Formula

              Volume(ml) = Desired dose / Dose in hand * Quantity.

     Here

                Desired dose (D) = 18 mg.

                Dose in hand (H) = 2 mg.

                      Quantity (Q) = 1 ml

    As per above formula

            Volume(ml) of drug required = D/H * Q

                                                          = 18/2 * 1

                                                          = 9 * 1

                                                          = 9 ml.